The plate current in a triode can be written as
i base p = k(V base g + Vp/μ)^3/2.
Show that the mutual conductance is proportional to the cube root of the plate current.

Grade:10

1 Answers

Jitender Pal
9 years ago
Sol. I base p = K(V base g + V base p/μ)^3/2 …(1) Diff. the equation : di base p = K 3/2 (V bsae g + V base p/μ)^1/2 d V base g ⇒ di base p/dV base g = 3/2 K (V base g + V base 0/μ)^1/23 ⇒ g base m = 3/2 K (V base g + V base p/μ)^1/2 …(2) From (1) i base p = [3/2 K (V base g + V base p/μ)^1/2]3 * 8/K^2 27 ⇒ i base p = k’ (g base m)^3 ⇒ g base m ∝ 3√i base p